### v1

@@ 2.2 Intelligence Efficiency and Local-Cloud Hybrid Systems
The viability of local inference on edge devices has been systematically characterized through the lens of efficiency metrics. Saad-Falcon et al. (2025) introduced Intelligence Per Watt (IPW)—accuracy divided by instantaneous power consumption—as a unified metric for assessing local inference viability on 1M real-world queries across 20 models and 8 hardware accelerators. Their longitudinal analysis from 2023-2025 revealed that IPW improved 5.3× through compounding advances in both model architectures (3.1× gains) and hardware accelerators (1.7× gains), with local LM coverage increasing from 23.2% to 71.3%. Critically, they demonstrated that intelligent routing between local and cloud models achieves 64.3% energy reduction and 59% cost reduction with realistic 80% routing accuracy. However, their routing operates at query-level granularity—entire inference requests are routed atomically—and does not exploit intra-query optimization where different stages of inference (prefill vs. decode) can be distributed across heterogeneous devices. Narayan et al. (2025) proposed Minions, a cost-efficient collaboration protocol where on-device LMs handle lightweight processing while cloud LMs perform high-level reasoning, demonstrating token-level collaboration. Our work extends this to fine-grained task-level routing with principled scaling law characterization across full heterogeneous device portfolios.
@@ 2.3 Heterogeneous Computing and Cost-Aware Orchestration
Efficient orchestration across heterogeneous hardware has become essential for sustainable AI deployment. Asgar et al. (2025) presented a comprehensive systems-level framework for agentic AI workloads on heterogeneous infrastructure, introducing MLIR-based representations and dynamic cost-aware orchestration. Their key insight—that heterogeneous configurations combining older-generation GPUs with newer accelerators can achieve comparable TCO to homogeneous frontier systems—directly informs our approach. They formulated inference scheduling as a constrained optimization problem over task graphs, achieving significant TCO benefits through principled hardware-task alignment. However, their focus remained on agentic workloads with tool calls, memory access, and multi-turn interactions, not on characterizing fundamental scaling law properties for pure inference, and they lacked empirical analysis of energy-coverage trade-offs. Meng and others (2024) developed an end-to-end framework for customizable neural network compression and deployment targeting edge hardware, addressing the critical hardware-software co-design gap. Zhang and others (2025) explored efficient inference on integrated edge processors, demonstrating that LayerNorm and hardware-aware optimization enable deployment on heterogeneous processors—architectural insights applicable to our lightweight student models.
@@ 2.4 Energy-Efficient Edge Deployment and Real-World Constraints
Deploying machine learning on resource-constrained devices imposes hard energy and memory budgets. Kannan and others (2022) established TinyML as a practical framework for deploying models on microcontrollers with kilobyte-scale memory budgets, demonstrating feasibility of machine learning on ultra-constrained devices. Pau and Zhuang (2024) synthesized rapid deployment methodologies for edge devices, emphasizing the importance of hardware-aware co-design and the trade-offs between latency, energy, and accuracy that our framework systematically characterizes. Adelola and others (2021) evaluated neural network compression methods for object detection on embedded systems, finding that pruning and quantization combinations yield optimal results for resource-constrained deployment—principles complementary to our inference-time scaling approach. Chen and others (2024) surveyed efficient deep learning for mobile devices, identifying key challenges in simultaneous optimization of model size, latency, and energy—the exact multi-objective landscape our QEIL framework addresses. These works collectively establish that hardware constraints are fundamental to edge deployment, yet none characterize how these constraints interact with inference-time scaling laws.
@@ 2.5 Limitations of Training-Time Scaling and the Case for Inference-Time Optimization
While scaling laws have become foundational in deep learning, recent critical analyses question whether the ”bigger is always better” paradigm remains viable. Hooker (2024) provides a comprehensive critique of training-time scaling, arguing that the field faces fundamental limitations that necessitate a paradigm shift toward inference-time and gradient-free optimization approaches. Hooker identifies four critical limitations of the traditional scaling paradigm that directly motivate our heterogeneous inference framework:
@@
(1) Diminishing Returns and Energy Crisis: Hooker demonstrates that training cost has resulted in massive capital accumulation disparity, excluding academic researchers and smaller institutions. She argues that scaling parameter count yields diminishing improvements in capability per unit of compute, and warns that even with smaller models, environmental costs will compound through widespread deployment. How QEIL Addresses This: Our approach avoids retraining entirely by leveraging inference-time repeated sampling (Theorem 1), achieving 70% pass@k coverage without parameter scaling. Theorem 2 (Energy Scaling) quantifies energy-coverage trade-offs explicitly, and our heterogeneous orchestration (Theorem 5) reduces total inference energy by 47.7% compared to homogeneous approaches—demonstrating that modest-sized models with intelligent resource allocation outperform larger models from an energy-efficiency perspective. This validates Hooker’s call to ”shift compute budgets from training to inference” with concrete mathematical proof and empirical validation.
@@
(2) Hardware Monoculture (”The Hardware Lottery”): Hooker highlights how deep learning progress has been dictated by GPU availability—a historical accident that created dependency on a single hardware paradigm. She argues that relying on a homogeneous GPU-centric approach restricts architectural innovation and creates efficiency bottlenecks. How QEIL Addresses This: Our Theorem 5 (Device-Task Efficiency Compatibility) fundamentally rejects hardware monoculture by decomposing inference into heterogeneous tasks with explicit device affinity. We prove that compute-bound prefill (high arithmetic intensity $I\approx 2L/3$) optimally maps to frequency-optimized GPUs, while memory-bound decode ($I\approx 1$, KV-cache bottleneck) maps to bandwidth-optimized NPUs. This heterogeneous composition achieves a 4.82× improvement in Intelligence Per Watt—a direct response to Hooker’s concern that homogeneous infrastructure represents a fundamental efficiency bottleneck. By introducing CPU, GPU, and NPU coordination, QEIL breaks the ”GPU lottery” and demonstrates that diverse hardware ecosystems drive efficiency gains impossible with single-device approaches.
@@
(3) Predictability and Statistical Uncertainty: Hooker critiques existing scaling laws for being ”surprisingly lacking in accuracy” when predicting downstream capabilities and performance. She notes that power-law extrapolations often fail when applied beyond training data, and that ”scaling laws cannot predict everything.” How QEIL Addresses This: Our five architecture-agnostic theorems are designed specifically for inference-time properties and provide predictive frameworks with demonstrable accuracy. Theorem 1 (Coverage Scaling) establishes that $C(S,N,T)=1-\exp(-\alpha(N)\cdot N^{\beta}\cdot S^{\beta}\cdot T^{\delta})$ holds empirically with universal exponent $\beta\approx 0.7$ across diverse model families (GPT-2, Llama, Qwen). Unlike training scaling laws that struggle to predict capability emergence, our inference laws explicitly characterize the relationship between sample budget and coverage—a directly observable, measurable quantity. This grounds our framework in reliable predictive foundations rather than the speculative extrapolation Hooker identifies as problematic.
@@
Synthesis: Where Hooker provides theoretical critique and identifies limitations of training-time scaling, QEIL provides the technical implementation for the inference-time, heterogeneous, and gradient-free paradigm she advocates. Our five theorems and hardware-task orchestration directly address each limitation: we replace parameter scaling with sample scaling, replace homogeneous hardware with heterogeneous routing, replace speculative extrapolation with empirically-grounded inference laws, and replace gradient-based optimization with hardware-aware scheduling. The result is a concrete system that achieves the efficiency gains and democratized accessibility Hooker argues are necessary for sustainable AI progress.
2.6 Reinforcement Learning Scaling Laws and Inference-Time Reasoning
While training-time scaling laws focus on pre-training efficiency, recent work has characterized how reinforcement learning (RL) scales with compute. Khatri et al. (2025) conducted the first large-scale empirical study of RL compute scaling, analyzing 400,000+ GPU-hours of RL training and deriving predictive scaling laws for RL performance curves with respect to compute allocation. Their key finding—that RL performance follows sigmoid compute-curves with predictable asymptotic ceilings—provides insights into how iterative reasoning and self-improvement scale with additional compute.
@@ 2.6 Reinforcement Learning Scaling Laws and Inference-Time Reasoning
From a scientific perspective, our inference-time approach offers advantages over RL for edge deployment: (1) No retraining overhead: RL requires backpropagation and model updates, making it infeasible on edge devices with limited memory and compute. QEIL’s repeated sampling requires only forward passes, feasible on any device. (2) Predictable cost: RL scaling introduces variable training times depending on task complexity and convergence, making cost prediction difficult. QEIL’s Theorem 3 (Latency Scaling) provides deterministic latency estimates as a function of samples and hardware. (3) Hardware flexibility: RL training typically requires GPUs. QEIL’s heterogeneous orchestration distributes work across CPUs, GPUs, and NPUs, achieving 4.82× better efficiency per unit power. (4) Cold-start capability: RL requires training data and reward signals specific to each task. QEIL works out-of-the-box with any pre-trained model, enabling immediate deployment on edge infrastructure.
While RL scaling laws remain important for understanding model improvement during training, QEIL demonstrates that inference-time scaling combined with heterogeneous orchestration provides a more practical and efficient path to improved performance on edge devices, achieving 70% pass@k coverage with 47.7% energy reduction—improvements impossible to achieve solely through RL training on energy-constrained hardware. The complementary insights suggest that optimal deployment strategy combines compute-optimal training (informed by Khatri’s scaling laws) with compute-optimal inference (informed by QEIL’s framework), where training produces efficient base models and inference-time sampling provides rapid capability scaling without retraining.
@@ 2.7 Distributed Inference and Disaggregated Processing
Disaggregating inference into distinct stages enables heterogeneous hardware utilization. Athiwaratkun et al. (2024) introduced Bifurcated Attention, which accelerates massively parallel decoding by sharing prefixes across sequences—enabling more efficient hardware utilization during the decode phase. Kwon et al. (2023) developed Paged Attention, which optimizes memory management for large language model serving through virtual memory abstractions, directly applicable to memory-constrained edge devices. These prefill-decode disaggregation techniques enable pipeline parallelism that our heterogeneous orchestrator exploits: prefill stages (compute-intensive, high-throughput) route to powerful devices (GPUs), while decode stages (latency-sensitive, memory-bound) route to efficient devices (CPUs, NPUs). Chen et al. (2024) analyzed scaling laws for compound inference systems combining multiple LLM calls, demonstrating that performance improvement follows predictable patterns as system complexity increases—lending theoretical support to our inference-time scaling framework.
@@ 2.8 Sparse Models and Mixture of Experts
2.9 Scaling Laws and Training-Time Compute Efficiency
Fundamental scaling laws characterizing model performance as a function of training compute have been extensively characterized. Hestness et al. (2017) established empirically that deep learning scaling follows predictable power laws, with loss scaling as $L(N)=\alpha N^{-\beta}$ across multiple orders of magnitude of model size $N$ and dataset size $D$. Hoffmann et al. (2022) and Kaplan et al. (2020) refined these laws, determining compute-optimal allocation between model size and training tokens. Shao et al. (2024) extended scaling law analysis to retrieval-augmented systems, observing smooth scaling with datastore size. However, these scaling laws characterize training-time compute, not inference-time behavior, and they do not address the heterogeneous hardware constraints that dominate edge deployment. Our work complements this literature by establishing inference-time scaling laws analogous to training-time laws, with explicit dependence on sample budget, quantization precision, and device-specific efficiency factors.

### v2

@@ 2.2 Intelligence Efficiency and Local-Cloud Hybrid Systems
The viability of local inference on edge devices has been systematically characterized through the lens of efficiency metrics. Saad-Falcon et al. (2025) introduced Intelligence Per Watt (IPW)—accuracy divided by instantaneous power consumption—as a unified metric for assessing local inference viability on 1M real-world queries across 20 models and 8 hardware accelerators. Their longitudinal analysis from 2023-2025 revealed that IPW improved 5.3× through compounding advances in both model architectures (3.1× gains) and hardware accelerators (1.7× gains), with local LM coverage increasing from 23.2% to 71.3%. Critically, they found that intelligent routing between local and cloud models achieves 64.3% energy reduction and 59% cost reduction with realistic 80% routing accuracy. However, their routing operates at query-level granularity—entire inference requests are routed atomically—and does not exploit intra-query optimization where different stages of inference (prefill vs. decode) can be distributed across heterogeneous devices. Narayan et al. (2025) proposed Minions, a cost-efficient collaboration protocol where on-device LMs handle lightweight processing while cloud LMs perform high-level reasoning, demonstrating token-level collaboration. Our work extends this to fine-grained task-level routing with principled scaling characterization across full heterogeneous device portfolios.
@@ 2.3 Heterogeneous Computing and Cost-Aware Orchestration
Efficient orchestration across heterogeneous hardware has become essential for sustainable AI deployment. Asgar et al. (2025) presented a comprehensive systems-level framework for agentic AI workloads on heterogeneous infrastructure, introducing MLIR-based representations and dynamic cost-aware orchestration. Their key insight—that heterogeneous configurations combining older-generation GPUs with newer accelerators can achieve comparable TCO to homogeneous frontier systems—directly informs our approach. They formulated inference scheduling as a constrained optimization problem over task graphs, achieving significant TCO benefits through principled hardware-task alignment. However, their focus remained on agentic workloads with tool calls, memory access, and multi-turn interactions, not on characterizing fundamental scaling properties for pure inference, and they lacked empirical analysis of energy-coverage trade-offs. Meng and others (2024) developed an end-to-end framework for customizable neural network compression and deployment targeting edge hardware, addressing the critical hardware-software co-design gap. Zhang and others (2025) explored efficient inference on integrated edge processors, demonstrating that LayerNorm and hardware-aware optimization enable deployment on heterogeneous processors—architectural insights applicable to our lightweight student models.
@@ 2.4 Energy-Efficient Edge Deployment and Real-World Constraints
Deploying machine learning on resource-constrained devices imposes hard energy and memory budgets. Kannan and others (2022) established TinyML as a practical framework for deploying models on microcontrollers with kilobyte-scale memory budgets, demonstrating feasibility of machine learning on ultra-constrained devices. Pau and Zhuang (2024) synthesized rapid deployment methodologies for edge devices, emphasizing the importance of hardware-aware co-design and the trade-offs between latency, energy, and accuracy that our framework systematically characterizes. Adelola and others (2021) evaluated neural network compression methods for object detection on embedded systems, finding that pruning and quantization combinations yield optimal results for resource-constrained deployment—principles complementary to our inference-time scaling approach. Chen and others (2024) surveyed efficient deep learning for mobile devices, identifying key challenges in simultaneous optimization of model size, latency, and energy—the exact multi-objective landscape our QEIL framework addresses. These works collectively establish that hardware constraints are fundamental to edge deployment, yet none characterize how these constraints interact with inference-time scaling relationships.
2.5 AI Safety, Reliability, and Fault-Tolerant Systems
The deployment of AI systems on edge devices introduces critical safety and reliability challenges that have received increasing attention. Amodei et al. (2016) established foundational concerns in AI safety, emphasizing the importance of safe exploration and robustness to distributional shift—concerns directly applicable to edge deployment where inference occurs in uncontrolled environments. Hendrycks et al. (2021) introduced benchmarks for measuring robustness to natural adversarial examples, demonstrating that model predictions can be highly sensitive to input perturbations—a vulnerability that edge systems must address through input validation and anomaly detection.
From a systems reliability perspective, Patterson et al. (2002) established that hardware failures are inevitable at scale, motivating fault-tolerant design principles. For edge AI specifically, this translates to graceful degradation when individual accelerators fail—a capability our orchestration framework implements through automatic workload redistribution. Avizienis et al. (2004) formalized fundamental concepts of dependable computing, including fault tolerance, error detection, and system recovery, which inform our safety-first design philosophy.
Thermal management is particularly critical for edge deployment. Pedram and Nazarian (2006) demonstrated that thermal throttling significantly impacts processor performance, establishing the need for thermal-aware workload scheduling. For mobile and edge devices specifically, Pathak et al. (2012) showed that energy consumption and thermal behavior are tightly coupled, and that aggressive computation can lead to thermal runaway. Our framework addresses this through proactive thermal monitoring and workload throttling before devices reach critical temperatures.
Adversarial robustness in deployed systems has been extensively studied. Goodfellow et al. (2015) introduced adversarial examples and demonstrated their transferability across models. For production systems, Carlini and Wagner (2017) showed that many proposed defenses are ineffective against adaptive attacks. Our approach implements defense-in-depth: input validation to reject malformed requests, output sanity checking to detect anomalous model behavior, and rate limiting to prevent denial-of-service attacks on edge resources.
Recent work on responsible AI deployment emphasizes the importance of “safety margins” in system design. Amodei and Clark (2016) argued that AI systems should be designed to fail gracefully rather than catastrophically. Our “safety-first, capability-second” philosophy directly implements this principle: we accept modest efficiency reductions (3-5%) to guarantee that inference never damages hardware, never produces unbounded resource consumption, and always maintains user-controllable behavior. This approach contrasts with pure performance optimization that may push devices beyond safe operating limits.
@@ 2.6 Limitations of Training-Time Scaling and the Case for Inference-Time Optimization
While scaling relationships have become foundational in deep learning, recent critical analyses question whether the ”bigger is always better” paradigm remains viable. Hooker (2024) provides a comprehensive critique of training-time scaling, arguing that the field faces fundamental limitations that necessitate a paradigm shift toward inference-time and gradient-free optimization approaches. Hooker identifies four critical limitations of the traditional scaling paradigm that directly motivate our heterogeneous inference framework:
@@
(1) Diminishing Returns and Energy Crisis: Hooker demonstrates that training cost has resulted in massive capital accumulation disparity, excluding academic researchers and smaller institutions. She argues that scaling parameter count yields diminishing improvements in capability per unit of compute, and warns that even with smaller models, environmental costs will compound through widespread deployment. How QEIL Addresses This: Our approach avoids retraining entirely by leveraging inference-time repeated sampling (Scaling Formalism 1), achieving 70% pass@k coverage without parameter scaling. Scaling Formalism 2 (Energy Scaling) quantifies energy-coverage trade-offs explicitly, and our heterogeneous orchestration (Scaling Formalism 5) reduces total inference energy by 47.7% compared to homogeneous approaches—suggesting that modest-sized models with intelligent resource allocation can outperform larger models from an energy-efficiency perspective. This validates Hooker’s call to “shift compute budgets from training to inference” with concrete empirical validation.
@@
(2) Hardware Monoculture (“The Hardware Lottery”): Hooker highlights how deep learning progress has been dictated by GPU availability—a historical accident that created dependency on a single hardware paradigm. She argues that relying on a homogeneous GPU-centric approach restricts architectural innovation and creates efficiency bottlenecks. How QEIL Addresses This: Our Scaling Formalism 5 (Device-Task Efficiency Compatibility) fundamentally rejects hardware monoculture by decomposing inference into heterogeneous tasks with explicit device affinity. We observe that compute-bound prefill (high arithmetic intensity $I\approx 2L/3$) optimally maps to frequency-optimized GPUs, while memory-bound decode ($I\approx 1$, KV-cache bottleneck) maps to bandwidth-optimized NPUs. This heterogeneous composition achieves a 4.82× improvement in Intelligence Per Watt—a direct response to Hooker’s concern that homogeneous infrastructure represents a fundamental efficiency bottleneck. By introducing CPU, GPU, and NPU coordination, QEIL breaks the “GPU lottery” and suggests that diverse hardware ecosystems drive efficiency gains impossible with single-device approaches.
@@
(3) Predictability and Statistical Uncertainty: Hooker critiques existing scaling relationships for being ”surprisingly lacking in accuracy” when predicting downstream capabilities and performance. She notes that power-law extrapolations often fail when applied beyond training data, and that ”scaling relationships cannot predict everything.” How QEIL Addresses This: Our five empirically validated scaling formalisms are designed specifically for inference-time properties and provide predictive frameworks with demonstrable accuracy. Scaling Formalism 1 (Coverage Scaling) establishes that $C(S,N,T)=1-\exp(-\alpha(N)\cdot N^{\beta_{N}}\cdot S^{\beta_{S}}\cdot T^{\delta})$ holds empirically with exponents $\beta_{N}\approx 0.7$ and $\beta_{S}\approx 0.7$ across diverse transformer model families (GPT-2, Llama, Qwen). Unlike training scaling relationships that struggle to predict capability emergence, our inference formalisms explicitly characterize the relationship between sample budget and coverage—a directly observable, measurable quantity. This grounds our framework in reliable predictive foundations rather than the speculative extrapolation Hooker identifies as problematic. Note: We use separate exponents $\beta_{N}$ for model size and $\beta_{S}$ for sample count to allow independent characterization of their respective contributions, though empirically they are approximately equal ($\beta_{N}\approx\beta_{S}\approx 0.7$) on our tested models.
@@
Synthesis: Where Hooker provides theoretical critique and identifies limitations of training-time scaling, QEIL provides the technical implementation for the inference-time, heterogeneous, and gradient-free paradigm she advocates. Our five scaling formalisms and hardware-task orchestration directly address each limitation: we replace parameter scaling with sample scaling, replace homogeneous hardware with heterogeneous routing, replace speculative extrapolation with empirically-grounded inference formalisms, and replace gradient-based optimization with hardware-aware scheduling. The result is a concrete system that achieves the efficiency gains and democratized accessibility Hooker argues are necessary for sustainable AI progress.
2.7 Reinforcement Learning Scaling and Inference-Time Reasoning
While training-time scaling focuses on pre-training efficiency, recent work has characterized how reinforcement learning (RL) scales with compute. Khatri et al. (2025) conducted the first large-scale empirical study of RL compute scaling, analyzing 400,000+ GPU-hours of RL training and deriving predictive relationships for RL performance curves with respect to compute allocation. Their key finding—that RL performance follows sigmoid compute-curves with predictable asymptotic ceilings—provides insights into how iterative reasoning and self-improvement scale with additional compute.
@@ 2.7 Reinforcement Learning Scaling and Inference-Time Reasoning
From a scientific perspective, our inference-time approach offers advantages over RL for edge deployment: (1) No retraining overhead: RL requires backpropagation and model updates, making it infeasible on edge devices with limited memory and compute. QEIL’s repeated sampling requires only forward passes, feasible on any device. (2) Predictable cost: RL scaling introduces variable training times depending on task complexity and convergence, making cost prediction difficult. QEIL’s Scaling Formalism 3 (Latency Scaling) provides deterministic latency estimates as a function of samples and hardware. (3) Hardware flexibility: RL training typically requires GPUs. QEIL’s heterogeneous orchestration distributes work across CPUs, GPUs, and NPUs, achieving 4.82× better efficiency per unit power. (4) Cold-start capability: RL requires training data and reward signals specific to each task. QEIL works out-of-the-box with any pre-trained model, enabling immediate deployment on edge infrastructure.
While RL scaling relationships remain important for understanding model improvement during training, QEIL suggests that inference-time scaling combined with heterogeneous orchestration provides a more practical and efficient path to improved performance on edge devices, achieving 70% pass@k coverage with 47.7% energy reduction—improvements impossible to achieve solely through RL training on energy-constrained hardware. The complementary insights suggest that optimal deployment strategy combines compute-optimal training (informed by Khatri’s scaling relationships) with compute-optimal inference (informed by QEIL’s framework), where training produces efficient base models and inference-time sampling provides rapid capability scaling without retraining.
@@ 2.8 Distributed Inference and Disaggregated Processing
Disaggregating inference into distinct stages enables heterogeneous hardware utilization. Athiwaratkun et al. (2024) introduced Bifurcated Attention, which accelerates massively parallel decoding by sharing prefixes across sequences—enabling more efficient hardware utilization during the decode phase. Kwon et al. (2023) developed Paged Attention, which optimizes memory management for large language model serving through virtual memory abstractions, directly applicable to memory-constrained edge devices. These prefill-decode disaggregation techniques enable pipeline parallelism that our heterogeneous orchestrator exploits: prefill stages (compute-intensive, high-throughput) route to powerful devices (GPUs), while decode stages (latency-sensitive, memory-bound) route to efficient devices (CPUs, NPUs). Chen et al. (2024) analyzed scaling relationships for compound inference systems combining multiple LLM calls, demonstrating that performance improvement follows predictable patterns as system complexity increases—lending theoretical support to our inference-time scaling framework.
@@ 2.9 Sparse Models and Mixture of Experts
2.10 Scaling Relationships and Training-Time Compute Efficiency
Fundamental scaling relationships characterizing model performance as a function of training compute have been extensively characterized. Hestness et al. (2017) established empirically that deep learning scaling follows predictable power laws, with loss scaling as $L(N)=\alpha N^{-\beta}$ across multiple orders of magnitude of model size $N$ and dataset size $D$. Hoffmann et al. (2022) and Kaplan et al. (2020) refined these relationships, determining compute-optimal allocation between model size and training tokens. Shao et al. (2024) extended scaling relationship analysis to retrieval-augmented systems, observing smooth scaling with datastore size. However, these scaling relationships characterize training-time compute, not inference-time behavior, and they do not address the heterogeneous hardware constraints that dominate edge deployment. Our work complements this literature by establishing inference-time scaling formalisms analogous to training-time scaling relationships, with explicit dependence on sample budget, quantization precision, and device-specific efficiency factors.